OpenStudy (anonymous):

factor completely 4x^2-16

hartnn (hartnn):

can you factor out 4 first ?

hartnn (hartnn):

\(4x^2= 4 \times x^2 \\ 16 = 4\times 4\)

OpenStudy (anonymous):

4(x^2-4) ?

OpenStudy (solomonzelman):

Yes

hartnn (hartnn):

now since we know that \(4 = 2^2\) we can use the relation \(\large a^2-b^2 = (a+b)(a-b)\) with a being 'x' and b being 2

OpenStudy (anonymous):

@SolomonZelman no 4(x^2-4) is not the answer. I got it wrong

hartnn (hartnn):

because we have still not completed the solution yet

OpenStudy (solomonzelman):

Yeah, SAD !